please help with these two example, i want to double check my work. thanks

1.

sum:=0

for i= 1 to 5
sum:=sum + i
end- for


2.

sum:=0

for i= 4 to 12
sum:=sum + i
end - for
Answer #1

1)
sum = 1+2+3+4+5 = 15
Answer: 15

2)
sum = 4+5+6+7+8+9+10+11+12 = 72
Answer: 72